OpenStudy (anonymous):

PLEASEEE ITS DUE TOMORROW :((( IT SAYS MY ANSWER IS WRONG Enter the equation of the line meeting the given conditions. Please put the equation in standard form. Containing E(4, 3) and F(6, 1)

OpenStudy (anonymous):

 is what i got

OpenStudy (anonymous):

(y-3) = (1-3)/6-4) (x-4) y-3 =-(x-4) or y-3=-x+4 y=-x+7 or x+y=7 ur answer is correct whats the proble then???

OpenStudy (anonymous):

is it in standard form? i know i really dont know whats wrong... its an online course

OpenStudy (anonymous):

its general form

OpenStudy (hba):

Well,This is the formula we have to use... \[\huge\ (x-x_1)=\frac{ y_2-y_1 }{ x_2-x_1 }(x-x_1)\] Where, E(4, 3)=(x_1,y_1) and F(6, 1)=(x_2,y_2)

OpenStudy (anonymous):

ohhh it says standard form... and okay so then how would it be solved?